OpenStudy (anonymous):

If you have a scale of 1.5 cm = 2 ft on a scaled drawing of a truck, what is the actual length if the truck measures 11.25 cm on the drawing?

OpenStudy (anonymous):

I will help you if you help me

OpenStudy (daniellelovee):

do you have the drawing?

OpenStudy (anonymous):

no drawing was provided, and ok shady

OpenStudy (daniellelovee):

oh well then the best I can do is try to help (but I can't promise anything)

OpenStudy (daniellelovee):

I believe this is cross multiplication are you familiar with that term?

OpenStudy (anonymous):

yeah like this 3/4= 18/x

OpenStudy (daniellelovee):

yes exactly

OpenStudy (anonymous):

so 1.5c/11.25cm=2ft/x ?

OpenStudy (daniellelovee):

no 1.5/2=11.25/x

OpenStudy (anonymous):

ah,ok, well I was close :P

OpenStudy (daniellelovee):

lol yeah XD good try and btw I got 15ft and that problem is missing information such as how am I supposed to find the length when I only have a scale o.O? and a scale does not determine width nor area which are the components of the length formula